# Rich and Poor Texture Contrast: A Simple yet Effective Approach for AI-generated Image Detection

### 10 Dec 2023 v2

@@ 4.2 Detection Effectiveness
Table 2 shows the detection accuracy comparison between our approach and baselines. Due to the space limitation, the average precision comparison can be found in the supplement. Our approach outperforms baselines with a clear margin in the average detection accuracy. Although our approach only trained over ProGAN-based fake images, it still works effectively for a wide variety of diffusion-based fake images. Our approach can also identify fake images generated by impressive commercial APIs like Midjourney, whose algorithm is unknown. In terms of each generative model, our approach achieves satisfactory results in most cases. The performance only degrades for fake images generated by CycleGAN and GauGAN. The fake images generated by CycleGAN differ from those generated from random noise. The inputs of CycleGAN are real images, and the outputs are stylized images. For instance, CycleGAN transfers a horse into a zebra. The performance degradation of our detector may result from distribution discrepancy. DIRE-D performs effectively for most diffusion-based generative models. However, its accuracy drops significantly for GAN-based fake images. The reason is that the performance of DIRE depends on their assumption that diffusion-generated fake images are more likely to be reconstructed by a pre-trained diffusion model. The reconstruction error of real images is close to that of GAN-generated fake images. As a result, DIRE-D cannot identify GAN-based fake images. In addition, DIRE-D is trained with fake images generated by ADM, a diffusion-based generative model. It is not surprising that DIRE-D achieves high detection accuracy over diffusion-generated fake images. We also train DIRE detector from scratch over ProGAN-based fake images, which is the same as other baselines. DIRE-G can identify most GAN-generated fake images but fails to generalize to diffusion-generated fake images. Among all baselines, the average detection accuracy of LNP is closest to ours. It is still inferior to ours by 5% over average detection accuracy.
@@ 4.3 Ablation Studies
Average 73.34 85.88 86.57
Degradation -15.98 -3.43 -2.74
Table 4: The impact of each component for detection accuracy. B. and C. denote boundary and contrast, respectively. The standard detection accuracy of our approach is 89.31. Each component is conducive to improving the detection accuracy.
@@
Since our rich/poor texture reconstructed image is made up of various patches, the boundary between different patches results in blocking artifact. We adopt high-pass filters across the boundaries to extract variation trends of patch diversity, which is conducive to the final detection accuracy. We try to remove blocking artifact by restricting the convolution range of high-pass filters. Specifically, high-pass filters do NOT cross the patch boundary and are only applied inside each patch. In this case named w/o Boundary, the detection accuracy of our detector slightly drops about 3.4%.
@@ 5 Conclusions
ProGAN 100.00 99.36 100.00 99.99 99.67 99.83 95.19 52.75 99.81 100.00
StyleGan 90.17 78.02 85.20 87.05 91.75 91.08 83.03 51.31 84.93 92.77
BigGAN 71.17 81.97 77.40 67.33 77.75 85.62 70.12 49.70 95.08 95.80
CycleGAN 87.62 78.77 87.00 86.07 84.10 86.94 74.19 49.58 98.33 70.17
StarGAN 94.60 94.62 97.00 95.05 99.92 99.27 95.47 46.72 95.75 99.97
GauGAN 81.42 80.57 77.00 69.35 75.39 78.46 67.79 51.23 99.47 71.58
Stylegan2 86.91 66.19 83.30 87.28 94.64 85.32 75.31 51.72 74.96 89.55
whichfaceisreal 91.65 50.75 66.80 86.80 70.85 55.70 58.05 53.30 86.90 85.80
ADM 60.39 63.42 49.00 58.61 84.73 67.15 75.78 98.25 66.87 82.17
Glide 58.07 54.13 57.20 54.50 80.52 66.11 71.75 92.42 62.46 83.79
Midjourney 51.39 45.87 52.20 50.02 65.55 65.35 58.01 89.45 56.13 90.12
SDv1.4 50.57 38.79 51.00 51.70 85.55 63.02 49.74 91.24 63.66 95.38
SDv1.5 50.53 39.21 51.40 52.16 85.67 63.67 49.83 91.63 63.49 95.30
VQDM 56.46 77.80 55.10 52.86 74.46 72.99 53.68 91.90 85.31 88.91
wukong 51.03 40.30 51.70 50.76 82.06 59.55 54.46 90.90 70.93 91.07
DALLE2 50.45 34.70 52.80 49.25 88.75 65.45 66.48 92.45 50.75 96.60
Average 70.78 64.03 68.38 68.67 83.84 75.34 68.68 71.53 78.43 89.31
@@
ProGAN 100.00 99.99 100.00 100.00 99.99 100.00 99.08 58.79 100.00 100.00
StyleGan 99.83 88.98 99.50 99.23 98.60 98.31 91.74 56.68 97.56 98.96
BigGAN 85.99 93.62 90.70 81.79 84.32 92.93 75.25 46.91 99.27 99.42
CycleGAN 94.94 84.78 95.50 95.33 92.83 95.01 80.56 50.03 99.80 85.26
@@
GauGAN 90.82 82.84 88.30 84.99 78.85 95.43 72.15 47.34 99.98 81.33
Stylegan2 99.48 82.54 99.60 99.11 99.59 97.89 88.29 58.03 97.90 97.74
whichfaceisreal 99.85 55.85 93.30 95.21 91.45 57.99 60.13 59.02 96.73 95.26
ADM 75.67 61.77 94.10 73.11 94.20 72.95 85.84 99.79 86.81 93.40
Glide 72.28 52.92 77.50 66.76 88.86 80.42 78.35 99.54 83.81 94.04
Midjourney 66.24 46.09 70.00 56.82 76.86 71.86 61.86 97.32 74.00 96.48
SDv1.4 61.20 37.83 65.40 59.83 94.31 62.37 49.87 98.61 86.14 99.06
SDv1.5 61.56 37.76 65.70 60.37 93.92 62.85 49.52 98.83 85.84 99.06
VQDM 68.83 85.10 75.60 61.13 87.35 77.47 54.57 98.98 96.53 96.26
wukong 57.34 39.58 64.60 55.62 92.38 62.48 55.38 98.37 91.07 97.54
DALLE2 53.51 38.20 68.12 49.82 96.14 82.55 74.48 99.71 63.04 99.56
Average 80.41 67.96 84.23 77.40 91.85 81.91 73.53 75.54 91.12 95.72
@@
ProGAN 99.96 84.40 99.64 99.94 71.16 55.70 98.92 51.55 99.34 97.84
StyleGan 75.00 72.30 70.45 77.74 56.02 55.90 75.93 50.89 79.74 82.49
BigGAN 62.05 62.90 63.50 62.28 51.20 51.35 67.85 51.08 88.22 65.25
CycleGAN 83.16 71.35 80.96 86.98 57.27 57.15 74.68 50.53 96.14 71.99
StarGAN 79.26 83.79 92.02 88.72 50.75 51.68 83.08 41.55 95.32 60.21
GauGAN 69.89 64.99 66.91 65.69 50.06 49.70 77.84 39.31 98.61 73.71
Stylegan2 71.29 73.34 60.19 78.87 58.81 55.99 68.99 51.36 69.41 82.71
whichfaceisreal 79.10 50.25 59.90 86.40 50.12 52.90 59.05 54.35 70.20 79.40
ADM 51.28 67.72 50.01 51.62 51.28 46.50 62.49 91.92 64.41 62.64
Glide 51.92 66.13 51.99 51.59 50.97 45.09 73.11 91.38 64.14 68.01
Midjourney 50.90 56.61 50.85 49.85 51.60 62.32 58.90 90.78 55.43 57.87
SDv1.4 49.82 52.95 50.16 49.44 52.66 53.87 48.40 92.53 56.48 75.00
SDv1.5 49.90 52.71 50.13 49.56 52.31 53.62 48.48 92.56 56.19 74.87
VQDM 51.24 76.51 51.61 50.98 50.98 44.06 54.82 92.47 78.89 64.94
wukong 50.01 52.60 50.50 49.62 51.55 52.49 50.88 92.14 62.75 67.91
DALLE2 49.70 82.70 50.00 48.20 50.50 36.40 60.40 89.89 50.35 70.35
Average 64.03 66.95 62.43 65.47 53.58 51.55 66.49 70.27 74.10 72.48
@@
ProGAN 100.00 95.45 100.00 100.00 92.64 68.65 99.94 50.95 99.98 99.73
StyleGan 98.60 79.65 98.62 98.24 75.36 75.48 94.12 53.59 94.90 91.12
BigGAN 88.36 69.81 91.39 83.14 65.83 54.21 78.07 50.21 97.65 69.41
CycleGAN 96.08 82.42 97.13 96.32 82.86 70.26 84.29 52.13 99.47 83.3
StarGAN 93.63 94.37 99.29 97.16 68.50 61.44 92.16 39.44 99.07 70.29
GauGAN 95.53 68.76 96.69 89.54 57.67 51.42 76.22 37.73 99.91 82.41
Stylegan2 98.08 82.48 98.26 98.42 85.29 77.57 92.02 54.50 94.65 91.03
whichfaceisreal 94.34 54.14 95.27 95.19 69.25 55.02 61.86 60.36 88.13 89.49
ADM 64.18 87.14 88.56 60.64 60.73 44.73 71.20 98.99 84.00 75.07
Glide 69.60 91.06 80.01 62.67 60.94 42.16 83.60 98.44 85.18 81.04
Midjourney 61.95 72.61 75.21 54.79 65.18 68.04 65.10 97.96 71.41 66.42
SDv1.4 55.52 66.99 64.10 52.56 68.26 56.03 45.14 99.15 75.23 86.47
SDv1.5 56.47 66.15 64.28 53.32 67.23 56.30 45.22 99.19 74.82 86.59
VQDM 64.82 91.26 76.46 56.98 57.90 40.91 58.94 99.39 93.58 78.54
wukong 54.78 65.54 63.78 51.40 63.69 54.32 50.50 99.23 83.69 80.26
DALLE2 47.03 98.58 54.97 43.52 49.61 36.37 66.37 99.51 58.58 95.59
Average 77.43 79.15 84.00 74.62 68.18 57.06 72.80 74.42 87.52 83.10
@@
ProGAN 88.00 60.85 50.00 88.30 71.50 81.46 68.01 49.74 95.83 99.92
StyleGan 64.46 61.81 50.00 67.69 69.19 71.32 66.52 51.74 72.63 90.37
BigGAN 52.02 52.70 50.00 52.20 61.60 58.23 52.08 50.68 73.00 72.35
CycleGAN 60.83 48.75 50.00 67.41 67.71 52.42 57.53 46.21 89.21 83.76
StarGAN 64.58 51.68 50.00 77.89 56.08 58.38 59.80 58.39 88.02 99.90
GauGAN 65.61 50.55 50.00 63.97 49.73 55.04 44.82 52.61 91.23 62.07
Stylegan2 65.69 63.50 50.00 69.40 74.46 69.93 66.79 60.05 68.66 89.00
whichfaceisreal 76.50 49.70 50.00 79.80 55.05 56.70 50.85 50.00 73.75 79.55
ADM 50.09 15.68 49.98 49.38 53.97 55.89 54.63 75.27 71.94 71.12
Glide 50.21 17.76 49.94 48.87 48.52 58.52 58.92 71.67 69.56 58.37
Midjourney 51.92 16.03 50.35 51.47 54.00 62.26 53.67 69.56 50.40 57.87
SDv1.4 51.42 15.18 49.93 50.98 55.25 59.24 50.80 74.48 51.17 81.39
SDv1.5 51.66 14.91 49.94 50.76 55.15 58.82 49.36 74.39 51.04 81.01
VQDM 50.24 17.01 49.96 49.48 46.14 57.96 52.56 76.06 81.45 75.30
wukong 50.31 16.66 49.97 49.92 57.58 56.65 52.36 69.62 54.64 78.74
DALLE2 48.00 20.60 49.95 47.30 45.30 60.95 58.75 65.75 51.40 73.40
Average 58.85 35.84 50.00 60.30 57.58 60.86 56.09 62.26 70.87 78.36
@@
ProGAN 99.27 81.83 50.69 98.94 92.36 97.21 77.93 50.18 99.35 100.00
StyleGan 90.80 72.00 47.53 90.25 93.60 91.79 80.44 56.19 90.72 99.17
BigGAN 61.36 57.93 51.18 61.15 73.61 64.29 52.73 52.25 84.65 72.34
CycleGAN 77.25 54.03 50.74 82.52 78.42 58.24 60.26 46.40 96.34 92.49
StarGAN 93.20 73.29 53.17 95.17 99.30 97.29 69.22 47.37 95.72 100.00
GauGAN 85.10 54.90 49.60 82.65 50.30 57.91 45.73 40.67 97.51 67.57
Stylegan2 91.65 69.83 45.63 91.46 96.17 94.44 77.45 70.22 87.12 99.17
whichfaceisreal 84.73 49.15 45.04 86.02 68.15 61.07 51.34 51.33 89.00 91.22
ADM 61.52 32.85 33.52 57.21 59.82 59.18 60.45 99.62 91.77 81.07
Glide 60.63 33.03 40.11 55.83 47.40 65.90 71.15 99.53 91.14 65.03
Midjourney 54.47 31.72 71.05 54.78 55.27 67.83 58.95 99.18 52.57 63.24
SDv1.4 55.50 31.56 39.14 53.80 58.72 62.72 51.93 99.71 62.25 91.01
SDv1.5 55.97 31.51 39.22 53.68 58.79 63.31 50.42 99.71 62.56 91.10
VQDM 59.25 33.23 39.68 55.79 47.85 62.63 57.93 99.54 95.71 84.03
wukong 51.57 31.76 39.94 50.39 61.90 61.21 52.31 99.49 72.80 87.03
DALLE2 55.21 34.84 45.38 50.55 47.53 67.48 69.84 99.58 72.99 85.78
Average 71.09 48.34 46.35 70.01 68.08 70.78 61.75 75.69 83.89 85.55
@@
ProGAN 99.95 85.02 99.94 99.90 84.67 95.46 85.85 51.84 98.65 99.01
StyleGan 83.32 80.35 83.15 84.84 76.85 85.26 72.79 50.75 71.99 90.38
BigGAN 68.03 71.90 74.00 67.42 57.30 63.90 57.35 51.28 76.92 63.00
CycleGAN 85.65 69.91 83.61 86.41 54.39 53.48 65.44 50.34 94.66 75.47
StarGAN 87.12 84.67 95.82 92.95 78.06 92.25 80.55 41.52 89.62 78.71
GauGAN 79.30 58.99 81.08 72.98 52.00 61.09 62.72 39.23 97.46 60.65
Stylegan2 84.04 79.79 88.43 87.49 86.67 86.09 63.20 51.04 62.11 91.99
whichfaceisreal 82.70 49.60 69.70 82.40 47.75 53.35 61.15 52.80 58.55 62.30
ADM 59.30 65.81 55.83 57.88 77.05 70.99 65.63 93.13 64.50 69.58
Glide 55.28 67.34 53.27 54.45 82.86 78.54 75.14 92.44 60.88 72.52
Midjourney 51.37 51.60 51.30 50.68 54.21 77.43 55.44 90.42 55.48 76.28
SDv1.4 51.23 49.10 48.70 52.87 63.47 62.93 47.04 92.87 54.56 78.85
SDv1.5 51.46 48.89 49.07 52.94 64.03 63.32 47.21 92.87 54.61 78.61
VQDM 55.56 67.43 54.67 54.35 64.82 63.28 57.98 93.43 76.47 70.53
wukong 50.62 46.03 49.53 51.43 62.64 60.09 51.63 92.83 58.59 74.23
DALLE2 49.30 75.55 51.35 49.10 79.75 80.20 74.85 90.60 49.95 72.00
Average 68.39 65.75 68.09 68.63 67.91 71.73 64.00 70.46 70.31 75.99
@@
ProGAN 100.00 95.28 100.00 100.00 93.18 99.20 95.01 51.10 99.92 99.98
StyleGan 98.97 88.29 98.69 98.70 88.32 96.50 84.15 52.58 92.84 97.59
BigGAN 79.50 80.07 83.19 79.82 56.50 69.38 59.17 51.23 91.98 64.41
CycleGAN 91.80 76.50 90.70 94.84 59.92 57.65 71.58 51.60 98.95 80.92
StarGAN 97.92 93.58 99.69 98.81 96.68 99.78 82.09 42.96 96.00 98.21
GauGAN 87.61 61.68 86.76 84.55 49.20 63.91 55.99 39.11 99.67 65.98
Stylegan2 99.00 88.60 99.60 99.01 95.04 97.35 77.42 54.18 90.33 98.08
whichfaceisreal 87.87 49.99 90.48 91.14 43.83 57.24 61.85 58.79 68.65 75.42
ADM 69.64 74.58 66.12 68.40 86.83 80.35 68.50 99.66 80.74 78.88
Glide 63.81 76.10 60.85 62.91 91.03 86.81 80.90 98.98 77.25 83.25
Midjourney 55.45 48.82 56.30 55.03 57.70 86.39 58.39 97.62 65.53 88.03
SDv1.4 56.58 45.05 52.16 58.99 71.34 71.72 46.01 99.39 67.54 90.80
SDv1.5 56.90 45.06 52.01 59.65 72.51 72.31 46.05 99.51 67.24 90.65
VQDM 64.46 76.80 63.89 62.06 70.96 70.94 59.59 99.87 90.23 78.05
wukong 54.09 38.48 51.90 55.64 71.29 67.87 52.51 99.47 74.59 83.75
DALLE2 49.84 91.22 53.86 47.65 93.06 88.19 81.81 99.60 54.28 90.81
Average 75.84 70.63 75.39 76.08 74.84 79.10 67.56 74.73 82.23 85.40

### 07 Mar 2024 v3

@@ 4.2 Detection Effectiveness
Table 2 shows the detection accuracy comparison between our approach and baselines. Due to the space limitation, the average precision comparison can be found in the supplement. Our approach outperforms baselines with a clear margin in the average detection accuracy. Although our approach only trained over ProGAN-based fake images, it still works effectively for a wide variety of diffusion-based fake images. Our approach can also identify fake images generated by impressive commercial APIs like Midjourney, whose algorithm is unknown. In terms of each generative model, our approach achieves satisfactory results in most cases. The performance only degrades for fake images generated by CycleGAN and GauGAN. The fake images generated by CycleGAN differ from those generated from random noise. The inputs of CycleGAN are real images, and the outputs are stylized images. For instance, CycleGAN transfers a horse into a zebra. The performance degradation of our detector may result from distribution discrepancy. DIRE-D performs effectively for most diffusion-based generative models. However, its accuracy drops significantly for GAN-based fake images. The reason is that the performance of DIRE depends on their assumption that diffusion-generated fake images are more likely to be reconstructed by a pre-trained diffusion model. The reconstruction error of real images is close to that of GAN-generated fake images. As a result, DIRE-D cannot identify GAN-based fake images. In addition, DIRE-D is trained with fake images generated by ADM, a diffusion-based generative model. It is not surprising that DIRE-D achieves high detection accuracy over diffusion-generated fake images. We also train DIRE detector from scratch over ProGAN-based fake images, which is the same as other baselines. DIRE-G can identify most GAN-generated fake images but fails to generalize to diffusion-generated fake images. Among all baselines, the average detection accuracy of LNP is closest to ours. It is still inferior to ours by 4% over average detection accuracy.
@@ 4.3 Ablation Studies
SDXL 62.92 97.95 95.7
Average 73.34 85.88 87.10
Degradation -17.13 -3.26 -2.74
Table 4: The impact of each component for detection accuracy. B. and C. denote boundary and contrast, respectively. The standard detection accuracy of our approach is 89.85. Each component is conducive to improving the detection accuracy.
@@
Since our rich/poor texture reconstructed image is made up of various patches, the boundary between different patches results in blocking artifact. We adopt high-pass filters across the boundaries to extract variation trends of patch diversity, which is conducive to the final detection accuracy. We try to remove blocking artifact by restricting the convolution range of high-pass filters. Specifically, high-pass filters do NOT cross the patch boundary and are only applied inside each patch. In this case named w/o Boundary, the detection accuracy of our detector slightly drops about 3.26%.
@@ 5 Conclusions
ProGAN 100.00 99.36 100.00 99.99 99.95 99.83 95.19 52.75 99.81 100.00
StyleGan 90.17 78.02 85.20 87.05 92.64 91.08 83.03 51.31 84.93 92.77
BigGAN 71.17 81.97 77.40 67.33 88.43 85.62 70.12 49.70 95.08 95.80
CycleGAN 87.62 78.77 87.00 86.07 79.07 86.94 74.19 49.58 98.33 70.17
StarGAN 94.60 94.62 97.00 95.05 100.00 99.27 95.47 46.72 95.75 99.97
GauGAN 81.42 80.57 77.00 69.35 79.17 78.46 67.79 51.23 99.47 71.58
Stylegan2 86.91 66.19 83.30 87.28 93.82 85.32 75.31 51.72 74.96 89.55
whichfaceisreal 91.65 50.75 66.80 86.80 50.00 55.70 58.05 53.30 86.90 85.80
ADM 60.39 63.42 49.00 58.61 83.91 67.15 75.78 98.25 66.87 82.17
Glide 58.07 54.13 57.20 54.50 83.50 66.11 71.75 92.42 62.46 83.79
Midjourney 51.39 45.87 52.20 50.02 69.55 65.35 58.01 89.45 56.13 90.12
SDv1.4 50.57 38.79 51.00 51.70 89.33 63.02 49.74 91.24 63.66 95.38
SDv1.5 50.53 39.21 51.40 52.16 88.81 63.67 49.83 91.63 63.49 95.30
VQDM 56.46 77.80 55.10 52.86 85.03 72.99 53.68 91.90 85.31 88.91
wukong 51.03 40.30 51.70 50.76 86.39 59.55 54.46 90.90 70.93 91.07
DALLE2 50.45 34.70 52.80 49.25 92.45 65.45 66.48 92.45 50.75 96.60
SDXL 53.03 51.23 55.60 64.53 87.75 71.30 55.35 91.28 50.73 98.43
Average 69.73 63.28 67.63 68.43 85.28 75.11 67.90 72.70 76.80 89.85
@@
ProGAN 100.00 99.99 100.00 100.00 100.00 100.00 99.08 58.79 100.00 100.00
StyleGan 99.83 88.98 99.50 99.23 99.27 98.31 91.74 56.68 97.56 98.96
BigGAN 85.99 93.62 90.70 81.79 94.54 92.93 75.25 46.91 99.27 99.42
CycleGAN 94.94 84.78 95.50 95.33 89.52 95.01 80.56 50.03 99.80 85.26
@@
GauGAN 90.82 82.84 88.30 84.99 84.54 95.43 72.15 47.34 99.98 81.33
Stylegan2 99.48 82.54 99.60 99.11 99.70 97.89 88.29 58.03 97.90 97.74
whichfaceisreal 99.85 55.85 93.30 95.21 42.75 57.99 60.13 59.02 96.73 95.26
ADM 75.67 61.77 94.10 73.11 93.37 72.95 85.84 99.79 86.81 93.40
Glide 72.28 52.92 77.50 66.76 92.76 80.42 78.35 99.54 83.81 94.04
Midjourney 66.24 46.09 70.00 56.82 86.92 71.86 61.86 97.32 74.00 96.48
SDv1.4 61.20 37.83 65.40 59.83 96.34 62.37 49.87 98.61 86.14 99.06
SDv1.5 61.56 37.76 65.70 60.37 96.00 62.85 49.52 98.83 85.84 99.06
VQDM 68.83 85.10 75.60 61.13 94.91 77.47 54.57 98.98 96.53 96.26
wukong 57.34 39.58 64.60 55.62 95.33 62.48 55.38 98.37 91.07 97.54
DALLE2 53.51 38.20 68.12 49.82 98.26 82.55 74.48 99.71 63.04 99.56
SDXL 72.62 49.45 79.41 68.22 87.75 80.03 53.97 99.10 67.59 99.89
Average 79.95 66.87 83.95 76.86 91.29 81.80 72.38 76.92 89.73 96.07
@@
ProGAN 99.96 84.40 99.64 99.94 67.80 55.70 98.92 51.55 99.34 97.84
StyleGan 75.00 72.30 70.45 77.74 55.42 55.90 75.93 50.89 79.74 82.49
BigGAN 62.05 62.90 63.50 62.28 51.73 51.35 67.85 51.08 88.22 65.25
CycleGAN 83.16 71.35 80.96 86.98 62.34 57.15 74.68 50.53 96.14 71.99
StarGAN 79.26 83.79 92.02 88.72 52.00 51.68 83.08 41.55 95.32 60.21
GauGAN 69.89 64.99 66.91 65.69 50.18 49.70 77.84 39.31 98.61 73.71
Stylegan2 71.29 73.34 60.19 78.87 58.32 55.99 68.99 51.36 69.41 82.71
whichfaceisreal 79.10 50.25 59.90 86.40 50.13 52.90 59.05 54.35 70.20 79.40
ADM 51.28 67.72 50.01 51.62 51.22 46.50 62.49 91.92 64.41 62.64
Glide 51.92 66.13 51.99 51.59 51.28 45.09 73.11 91.38 64.14 68.01
Midjourney 50.90 56.61 50.85 49.85 50.14 62.32 58.90 90.78 55.43 57.87
SDv1.4 49.82 52.95 50.16 49.44 51.84 53.87 48.40 92.53 56.48 75.00
SDv1.5 49.90 52.71 50.13 49.56 51.87 53.62 48.48 92.56 56.19 74.87
VQDM 51.24 76.51 51.61 50.98 51.06 44.06 54.82 92.47 78.89 64.94
wukong 50.01 52.60 50.50 49.62 50.91 52.49 50.88 92.14 62.75 67.91
DALLE2 49.70 82.70 50.00 48.20 50.70 36.40 60.40 89.89 50.35 70.35
SDXL 49.98 75.00 50.48 49.93 49.93 65.98 56.53 91.13 50.65 69.00
Average 63.20 67.43 61.72 64.55 53.35 52.39 65.90 71.50 72.72 72.01
@@
ProGAN 100.00 95.45 100.00 100.00 91.04 68.65 99.94 50.95 99.98 99.73
StyleGan 98.60 79.65 98.62 98.24 79.86 75.48 94.12 53.59 94.90 91.12
BigGAN 88.36 69.81 91.39 83.14 67.56 54.21 78.07 50.21 97.65 69.41
CycleGAN 96.08 82.42 97.13 96.32 90.39 70.26 84.29 52.13 99.47 83.3
StarGAN 93.63 94.37 99.29 97.16 76.68 61.44 92.16 39.44 99.07 70.29
GauGAN 95.53 68.76 96.69 89.54 60.98 51.42 76.22 37.73 99.91 82.41
Stylegan2 98.08 82.48 98.26 98.42 87.56 77.57 92.02 54.50 94.65 91.03
whichfaceisreal 94.34 54.14 95.27 95.19 69.81 55.02 61.86 60.36 88.13 89.49
ADM 64.18 87.14 88.56 60.64 61.90 44.73 71.20 98.99 84.00 75.07
Glide 69.60 91.06 80.01 62.67 69.27 42.16 83.60 98.44 85.18 81.04
Midjourney 61.95 72.61 75.21 54.79 63.94 68.04 65.10 97.96 71.41 66.42
SDv1.4 55.52 66.99 64.10 52.56 70.53 56.03 45.14 99.15 75.23 86.47
SDv1.5 56.47 66.15 64.28 53.32 70.51 56.30 45.22 99.19 74.82 86.59
VQDM 64.82 91.26 76.46 56.98 60.51 40.91 58.94 99.39 93.58 78.54
wukong 54.78 65.54 63.78 51.40 66.00 54.32 50.50 99.23 83.69 80.26
DALLE2 47.03 98.58 54.97 43.52 55.31 36.37 66.37 99.51 58.58 95.59
SDXL 75.26 94.45 80.73 43.63 50.82 73.81 60.10 99.15 64.97 86.12
Average 77.31 80.05 83.81 72.80 70.16 58.04 72.05 75.88 86.19 83.11
@@
ProGAN 88.00 60.85 50.00 88.30 85.15 81.46 68.01 49.74 95.83 99.92
StyleGan 64.46 61.81 50.00 67.69 80.34 71.32 66.52 51.74 72.63 90.37
BigGAN 52.02 52.70 50.00 52.20 70.95 58.23 52.08 50.68 73.00 72.35
CycleGAN 60.83 48.75 50.00 67.41 67.79 52.42 57.53 46.21 89.21 83.76
StarGAN 64.58 51.68 50.00 77.89 86.44 58.38 59.80 58.39 88.02 99.90
GauGAN 65.61 50.55 50.00 63.97 53.89 55.04 44.82 52.61 91.23 62.07
Stylegan2 65.69 63.50 50.00 69.40 89.16 69.93 66.79 60.05 68.66 89.00
whichfaceisreal 76.50 49.70 50.00 79.80 51.90 56.70 50.85 50.00 73.75 79.55
ADM 50.09 15.68 49.98 49.38 66.25 55.89 54.63 75.27 71.94 71.12
Glide 50.21 17.76 49.94 48.87 53.83 58.52 58.92 71.67 69.56 58.37
Midjourney 51.92 16.03 50.35 51.47 49.06 62.26 53.67 69.56 50.40 57.87
SDv1.4 51.42 15.18 49.93 50.98 48.28 59.24 50.80 74.48 51.17 81.39
SDv1.5 51.66 14.91 49.94 50.76 47.81 58.82 49.36 74.39 51.04 81.01
VQDM 50.24 17.01 49.96 49.48 51.86 57.96 52.56 76.06 81.45 75.30
wukong 50.31 16.66 49.97 49.92 51.15 56.65 52.36 69.62 54.64 78.74
DALLE2 48.00 20.60 49.95 47.30 62.90 60.95 58.75 65.75 51.40 73.40
SDXL 61.25 24.88 60.30 56.69 59.03 78.28 52.63 90.70 50.75 77.68
Average 58.99 35.19 50.61 60.09 63.28 61.89 55.89 63.94 69.69 78.34
@@
ProGAN 99.27 81.83 50.69 98.94 94.12 97.21 77.93 50.18 99.35 100.00
StyleGan 90.80 72.00 47.53 90.25 93.49 91.79 80.44 56.19 90.72 99.17
BigGAN 61.36 57.93 51.18 61.15 75.77 64.29 52.73 52.25 84.65 72.34
CycleGAN 77.25 54.03 50.74 82.52 71.89 58.24 60.26 46.40 96.34 92.49
StarGAN 93.20 73.29 53.17 95.17 99.50 97.29 69.22 47.37 95.72 100.00
GauGAN 85.10 54.90 49.60 82.65 52.67 57.91 45.73 40.67 97.51 67.57
Stylegan2 91.65 69.83 45.63 91.46 97.11 94.44 77.45 70.22 87.12 99.17
whichfaceisreal 84.73 49.15 45.04 86.02 79.90 61.07 51.34 51.33 89.00 91.22
ADM 61.52 32.85 33.52 57.21 76.01 59.18 60.45 99.62 91.77 81.07
Glide 60.63 33.03 40.11 55.83 58.44 65.90 71.15 99.53 91.14 65.03
Midjourney 54.47 31.72 71.05 54.78 48.75 67.83 58.95 99.18 52.57 63.24
SDv1.4 55.50 31.56 39.14 53.80 47.75 62.72 51.93 99.71 62.25 91.01
SDv1.5 55.97 31.51 39.22 53.68 47.63 63.31 50.42 99.71 62.56 91.10
VQDM 59.25 33.23 39.68 55.79 54.33 62.63 57.93 99.54 95.71 84.03
wukong 51.57 31.76 39.94 50.39 52.09 61.21 52.31 99.49 72.80 87.03
DALLE2 55.21 34.84 45.38 50.55 72.05 67.48 69.84 99.58 72.99 85.78
SDXL 75.95 33.46 72.47 55.45 58.05 85.90 52.00 99.05 60.41 85.01
Average 71.38 47.47 47.89 69.16 69.39 71.67 61.18 77.06 82.51 85.60
@@
ProGAN 99.95 85.02 99.94 99.90 76.56 95.46 85.85 51.84 98.65 99.01
StyleGan 83.32 80.35 83.15 84.84 68.36 85.26 72.79 50.75 71.99 90.38
BigGAN 68.03 71.90 74.00 67.42 53.78 63.90 57.35 51.28 76.92 63.00
CycleGAN 85.65 69.91 83.61 86.41 52.65 53.48 65.44 50.34 94.66 75.47
StarGAN 87.12 84.67 95.82 92.95 63.16 92.25 80.55 41.52 89.62 78.71
GauGAN 79.30 58.99 81.08 72.98 49.23 61.09 62.72 39.23 97.46 60.65
Stylegan2 84.04 79.79 88.43 87.49 76.75 86.09 63.20 51.04 62.11 91.99
whichfaceisreal 82.70 49.60 69.70 82.40 52.05 53.35 61.15 52.80 58.55 62.30
ADM 59.30 65.81 55.83 57.88 70.30 70.99 65.63 93.13 64.50 69.58
Glide 55.28 67.34 53.27 54.45 83.31 78.54 75.14 92.44 60.88 72.52
Midjourney 51.37 51.60 51.30 50.68 68.52 77.43 55.44 90.42 55.48 76.28
SDv1.4 51.23 49.10 48.70 52.87 64.31 62.93 47.04 92.87 54.56 78.85
SDv1.5 51.46 48.89 49.07 52.94 64.17 63.32 47.21 92.87 54.61 78.61
VQDM 55.56 67.43 54.67 54.35 57.05 63.28 57.98 93.43 76.47 70.53
wukong 50.62 46.03 49.53 51.43 60.68 60.09 51.63 92.83 58.59 74.23
DALLE2 49.30 75.55 51.35 49.10 85.90 80.20 74.85 90.60 49.95 72.00
SDXL 56.53 59.30 55.73 59.25 73.25 79.53 54.20 92.08 50.78 76.50
Average 67.69 65.37 67.36 68.08 65.88 72.19 63.42 71.73 69.16 75.92
@@
ProGAN 100.00 95.28 100.00 100.00 87.11 99.20 95.01 51.10 99.92 99.98
StyleGan 98.97 88.29 98.69 98.70 79.52 96.50 84.15 52.58 92.84 97.59
BigGAN 79.50 80.07 83.19 79.82 54.31 69.38 59.17 51.23 91.98 64.41
CycleGAN 91.80 76.50 90.70 94.84 52.52 57.65 71.58 51.60 98.95 80.92
StarGAN 97.92 93.58 99.69 98.81 88.55 99.78 82.09 42.96 96.00 98.21
GauGAN 87.61 61.68 86.76 84.55 47.07 63.91 55.99 39.11 99.67 65.98
Stylegan2 99.00 88.60 99.60 99.01 88.05 97.35 77.42 54.18 90.33 98.08
whichfaceisreal 87.87 49.99 90.48 91.14 49.75 57.24 61.85 58.79 68.65 75.42
ADM 69.64 74.58 66.12 68.40 79.53 80.35 68.50 99.66 80.74 78.88
Glide 63.81 76.10 60.85 62.91 92.53 86.81 80.90 98.98 77.25 83.25
Midjourney 55.45 48.82 56.30 55.03 75.00 86.39 58.39 97.62 65.53 88.03
SDv1.4 56.58 45.05 52.16 58.99 68.81 71.72 46.01 99.39 67.54 90.80
SDv1.5 56.90 45.06 52.01 59.65 68.91 72.31 46.05 99.51 67.24 90.65
VQDM 64.46 76.80 63.89 62.06 62.44 70.94 59.59 99.87 90.23 78.05
wukong 54.09 38.48 51.90 55.64 63.56 67.87 52.51 99.47 74.59 83.75
DALLE2 49.84 91.22 53.86 47.65 95.03 88.19 81.81 99.60 54.28 90.81
SDXL 77.49 65.50 67.66 56.84 75.09 86.39 54.02 99.25 59.07 92.44
Average 75.94 70.33 74.93 74.94 72.22 79.53 66.77 76.17 80.87 85.72